OpenStudy (anonymous):

(125/64) raised 2/3 simplify the expression thank you for all help :)

OpenStudy (anonymous):

raise to the two thirds power means a) take the cube root b) square

OpenStudy (anonymous):

\[\sqrt[3]{125}=5\] because \(5^2=125\) and \[\sqrt[3]{64}=4\] so you can start with \[(\frac{5}{4})^2\]

OpenStudy (anonymous):

25/16 ??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thaank you